# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v1

@@ 6.2 Generalization
The generalization aspect of multi-objective deep learning remains relatively underexplored compared to its convergence properties. Cortes et al. [99] analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations. Sukenik et al. [100] consider the generalization for a broader class of scalarization methods. They show that the generalization bounds for individual objectives extend to MOO with scalarization. Both studies rely on the Rademacher complexity of the hypothesis class to establish algorithm-independent generalization bounds, which are unaffected by the training process.
Another line of research investigates the Tchebycheff scalarization method, which focuses on the sample complexity required to achieve generalization error within $\epsilon$ of the optimal objective value. Formally, given a set of $m$ distributions $\{{\mathcal{D}}_{i}\}_{i=1}^{m}$ and a hypothesis class ${\mathcal{H}}$, the goal is to find a (possibly randomized) hypothesis $h$ such that
@@
where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss on ${\mathcal{D}}_{i}$ with hypothesis $h$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance. It is first formulated by Awasthi et al. [101] as an open problem in 2023 and has since been addressed by several works [102, 103]. Haghtalab et al. are the first to show the lower bound for sample complexity of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class ${\mathcal{H}}$. By using a boosting framework, Peng [102] gives an algorithm that achieves an upper bound for sample complexity of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal. In a concurrent work, Zhang et al. [103] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$, addressing the open problem raised by Awasthi et al. [101].
In the online setting, Liu et al. [104] provide an adaptive online mirror descent algorithm that achieves regret on the order of $O\left(\frac{mv}{\sqrt{K}}\right)$, where $K$ is the number of iterations. Using a plain online-to-batch conversion scheme, this algorithm leads to an $O\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$. It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
Moreover, MoDo [95] introduces a different perspective by leveraging algorithm stability to derive the first set of algorithm-dependent generalization error bounds of gradient balancing algorithms.
@@ 7.1 Reinforcement Learning
Reinforcement Learning (RL) [105] involves training agents to make sequential decisions by maximizing cumulative rewards. Multi-Objective Reinforcement Learning (MORL) extends single-objective RL by utilizing a vector-valued reward function, i.e., ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$. MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following training objective:
@@
Radius Algorithm (RA) [106] employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem. However, a uniform distribution of preferences does not always lead to a uniform distribution of solutions. To address this, various methods focus on improving the configuration of preference vectors. For instance, Optimal Linear Support (OLS) [107, 108] adds new preference vectors dynamically during each iteration. Moreover, Prediction-guided MORL (PGMORL) [109] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
Although Lu et al. [5] demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization. CAPQL [110] introduces a scalarization method that combines linear scalarization with an entropy term. PPA [111] incorporates a term to align objective values more closely with preference vectors. Several studies [112, 113, 114, 115, 116, 117] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto optimal policies. Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
Another research direction focuses on learning the entire Pareto set [118, 119, 120, 121, 122, 123]. In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and the preference vector ${\bm{\alpha}}$. Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
@@ 7.2 Bayesian Optimization
Bayesian Optimization (BO) [124] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model, typically a Gaussian Process, to approximate the objective function and guide the search. BO balances exploration (sampling uncertain regions) and exploitation (sampling promising regions) by optimizing an acquisition function like Expected Improvement (EI) [125] or Upper Confidence Bound (UCB) [126].
When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO). MOBO facilitates the exploration of trade-offs between different objectives and aims to approximate the Pareto set of optimal solutions.
In MOBO, acquisition functions must be designed to manage the balance between exploration and exploitation across multiple objectives. Based on this, MOBO methods can be classified into three main categories:
@@
    Scalarization-based Methods. These methods convert multi-objective problems into single-objective ones using scalarization functions discussed in Section 2.2 [127, 128]. Zhang and Golovin [129] propose using the hypervolume indicator for scalarization. Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [130], PAL [131, 132], and MOBO-RS [133] extend single-objective acquisitions to multi-objective optimization, leading to improved performance.
@@
    Information Theoretic-Based Methods. These methods aim to maximize information gain about the PF by selecting candidate points. PESMO [143] reduces uncertainty in the posterior over the Pareto optimal input set. Extensions like MESMO [144] and Pareto-Frontier Entropy Search [145] refine this by focusing on output space and the PF. Joint Entropy Search (JES)[146, 147] measures combined information gain across input and output spaces. USeMO[148] offers an alternative uncertainty measure based on hyper-rectangle volumes.
The solutions obtained by the aforementioned methods provide discrete approximations of the Pareto set. Lin et al. [149] propose learning a continuous approximation of Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1). By adjusting preferences, decision-makers can flexibly explore any trade-off region within the approximate Pareto set. Subsequent improvements have been proposed to leverage diffusion model [150] and stein variational gradient descent [151] for efficient solution generation, as well as incorporate data augmentation for Pareto set model building [152].
@@ 7.3 Computer Vision
The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has successfully applied in autonomous driving [153, 154]. Since the encoder in computer vision usually contains large parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in partial tasks. Hence, many methods are proposed to address the issue from the perspective of MOO, which use some scalarization methods to balance multiple losses [155, 156, 157, 158, 159]. A comprehensive survey on multi-task dense prediction is provided in [4].
Expect the dense prediction task, MOO is also useful in other computer vision tasks, such as point cloud [160, 161, 162], medical image denoising [163], and pose estimation [164].
@@ 7.4 Neural Architecture Search
Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently [165]. Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy. For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, latency, and energy consumption. These works are mainly based on gradient-based NAS methods (like DARTS [166]) and formulate it as a multi-objective optimization problem. Among them, [8, 167, 168, 169, 170] employ some scalarization methods to identify a single solution, where task and efficiency objectives are combined with fixed weights. However, altering the weights for objectives needs a complete rerun of the search, which is computationally intensive. Therefore, Sukthanker et al. [171] propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
@@ 7.5 Recommendation System
Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty and diversity, serendipity, popularity, and others [172]. These factors naturally frame the problem as a multi-objective optimization optimization. Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives [173, 174, 175]. In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model [176, 177, 178, 179]. Additionally, some research efforts aim to identify finite Pareto sets [180] and infinite Pareto sets [181, 182], enabling the provision of personalized recommender models that cater to varying user preferences.
@@ 7.6 Large Language Models (LLMs)
Due to the powerful transferability of LLMs, users can fine-tune LLMs to specific downstream tasks or scenarios. This approach separates fine-tuning on each task, causing extensive costs in training and difficulties in deployment. MFTCoder [183] proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks. CoBa [184] studies the multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects. MORLHF [185] and MODPO [186] train an LLM for every preference configuration by linearly combining multiple (implicit) reward models. To avoid the LLM retraining, some methods separately train multiple LLMs for each preference dimension and merge their parameters [187, 188] or output logits [189] to deal with different preference requirements in the inference. However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs. To improve the efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [190, 191, 192] or model parameters [54].
@@ 7.7 Miscellaneous
In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [193, 194], federated learning [195, 196, 197], long-tailed learning [198, 199, 200], continual learning [201], diffusion models [202, 203, 204, 205], neural combinatorial optimization [206, 119, 207, 208, 209], and GFlowNets [210, 211]. These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
@@ 8 Resources
8.1 Benchmark Datasets
In this section, we introduce some benchmark datasets for multi-objective deep learning. They are summarized in Table III and listed in detail as follows.
@@ 8.1 Benchmark Datasets
8.2 Software
In this section, we introduce two open-source libraries for gradient-based multi-objective deep learning: LibMTL [^1] https://github.com/median-research-group/LibMTL [225] and LibMOON [^2] https://github.com/xzhang2523/libmoon [226]. These two libraries provide a unified environment for users to implement or fairly evaluate different MOO methods. Moreover, they are built on PyTorch [227] and designed in a modular fashion, allowing users to flexibly develop novel methods or apply existing MOO methods to new application scenarios. LibMTL [225] mainly focuses on finding a single Pareto optimal solution and supports over $19$ methods introduced in Section 3. Meanwhile, it also supports $6$ benchmark datasets (i.e., NYUv2, Cityscapes, QM9, Office-31, Office-Home, and XTREME datasets) introduced in Section 8.1. LibMOON [226] mainly focuses on exploring the whole Pareto set. It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5). These methods address a variety of tasks, including classification, fairness, and distribution matching.
@@ 9.1 Theoretical Understanding
While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
@@
Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received much less attention. For instance, Chen et al. [95], as discussed in Section 6.2, provide the first algorithm-dependent generalization analysis for MOO. Extending this work to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
@@
For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets. Zhang et al. [228] established a generalization bound for hypervolume-based Pareto set learning. However, the effect of various network architectures (Section 5.1) on approximating Pareto sets and convergence rate remain unclear. Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
9.2 Reducing Gradient Balancing Costs
@@
While gradient balancing methods in Section 3.2 often show strong performance in various applications, they come with significant computational overhead. Despite the introduction of some practical speedup strategies (discussed in Section 3.2.3), existing approaches remain insufficiently efficient. A deeper understanding of the optimization differences between gradient balancing, linear scalarization, and loss balancing is crucial. This insight could facilitate the integration of gradient balancing with linear scalarization and loss balancing, reducing computational overhead significantly and enabling its application in large-scale training scenarios.
9.3 Dealing with Large Number of Objectives
@@
Some real-world problems involve handling a large number of objectives, which poses significant challenges for current Pareto set learning algorithms in Section 5. As the number of objectives grows, the preference vector space expands exponentially, making it challenging for existing random sampling-based techniques to effectively learn the mapping between preference vectors and solutions. Developing efficient sampling strategies for high-dimensional preference spaces remains a critical research problem. Additionally, exploring methods to reduce or merge objectives based on their properties could be a promising approach to minimize the total number of objectives.
9.4 Distributed Training
@@
Current MOO algorithms in deep learning are limited to single GPUs or machines, but scaling them to multi-GPU and distributed environments is increasingly important as models and datasets grow. This scaling introduces unique challenges compared to single-objective optimization. Gradient-balancing algorithms require efficient gradient distribution and communication across GPUs. Learning a set of Pareto optimal solutions involves synchronizing multiple models across nodes with minimal overhead. Additionally, when data for different objectives is distributed across devices and cannot be shared (e.g., due to privacy or memory constraints), collaborative computation of solutions or Pareto sets without direct data sharing becomes an important challenge.
9.5 Advancements in Large Language Models (LLMs)
@@
As discussed in Section 7.6, current research in MOO for LLMs largely focuses on the RLHF stage. Expanding MOO techniques to other stages in the LLM lifecycle, such as pre-training and inference, is a valuable direction for future research. Addressing these challenges could lead to models that are better aligned with user needs across their entire development processes. Additionally, user preferences are currently modeled as a preference vector. However, this representation may oversimplify more complex user preferences on LLMs. Exploring advanced methods to capture and represent these intricate preferences could enhance the effectiveness and personalization of LLMs.
9.6 Application in More Deep Learning Scenarios
@@
While MOO has already been utilized in various deep learning scenarios, as highlighted in Section 7, there remain numerous unexplored areas within this field. Multi-objective characteristics are inherently present in most deep learning problems, as methods are typically developed or models evaluated based on multiple criteria. Consequently, trade-offs often arise naturally. Leveraging MOO methods to effectively address these trade-offs presents a significant opportunity for further exploration.

### v2

@@ 6.2 Generalization
The generalization aspect of multi-objective deep learning remains relatively underexplored compared to its convergence properties. Cortes et al. [101] analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations. Sukenik et al. [102] consider a broader class of scalarization methods. They show that the generalization bounds for individual objectives extend to MOO with scalarization. Both studies rely on the Rademacher complexity of the hypothesis class to establish algorithm-independent generalization bounds, which are unaffected by the training process.
Another line of research investigates Tchebycheff scalarization, which focuses on the sample complexity required to achieve a generalization error within $\epsilon$ of the optimal objective value. Formally, given a set of $m$ distributions $\{{\mathcal{D}}_{i}\}_{i=1}^{m}$ and a hypothesis class ${\mathcal{H}}$, the goal is to find a (possibly randomized) hypothesis $h$ such that
@@
where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss of hypothesis $h$ on ${\mathcal{D}}_{i}$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance. It is first formulated by Awasthi et al. [103] as an open problem in 2023 and has since been addressed by several works [104, 105]. Haghtalab et al. are the first to show the sample complexity lower bound of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class ${\mathcal{H}}$. By using a boosting framework, Peng [104] gives an algorithm that achieves a sample complexity upper bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal. In a concurrent work, Zhang et al. [105] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$.
In the online setting, Liu et al. [106] provide an adaptive online mirror descent algorithm that achieves $\mathcal{O}\left(\frac{mv}{\sqrt{K}}\right)$ regret, where $K$ is the number of iterations. Using a plain online-to-batch conversion scheme, this algorithm leads to an $\mathcal{O}\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, which still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$ [105]. It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
MoDo [24] introduces a different perspective by leveraging algorithm stability to derive algorithm-dependent generalization error bounds of gradient balancing algorithms.
@@ 7.1 Reinforcement Learning
Reinforcement Learning (RL) [107] involves training agents to make sequential decisions by maximizing cumulative rewards. Multi-Objective Reinforcement Learning (MORL) [108] extends single-objective RL by utilizing a vector-valued reward function ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$. MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following training objective:
@@
Radius Algorithm (RA) [109] employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem. However, a uniform distribution of preferences does not always lead to a uniform distribution of solutions. To address this, various methods focus on improving the configuration of preference vectors. For instance, Optimal Linear Support (OLS) [110, 111] adds new preference vectors dynamically during each iteration. Prediction-guided MORL (PGMORL) [112] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
Although Lu et al. [5] demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization. CAPQL [113] introduces a scalarization method that combines linear scalarization with an entropy term. PPA [114] incorporates a term to align objective values more closely with preference vectors. Several studies [115, 116, 117, 118, 119, 120] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto-optimal policies. Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
Another research direction focuses on learning the entire Pareto set [121, 122, 123, 124, 125, 126]. In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and preference vector ${\bm{\alpha}}$. Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
@@ 7.2 Bayesian Optimization
Bayesian Optimization (BO) [127] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model. When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO) [128]. In MOBO, acquisition functions are designed to manage the balance between exploration and exploitation across multiple objectives. Based on this, MOBO methods can be classified into three main categories:
@@
    Scalarization-based Methods. These methods convert MOO problems into single-objective ones using scalarization functions discussed in Section 2.2 [128, 129]. Zhang and Golovin [130] propose using the hypervolume indicator for scalarization. Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [131], PAL [132, 133], and MOBO-RS [134] extend single-objective acquisitions to multi-objective optimization.
@@
    Information Theoretic-Based Methods. These methods aim to maximize information gain about the PF by selecting candidate points. PESMO [145] reduces uncertainty in the posterior over the Pareto optimal input set. Some extensions [146, 147, 148, 149] refine this by considering output space. USeMO[150] offers an alternative uncertainty measure based on hyper-rectangle volumes.
The solutions obtained by the aforementioned methods provide discrete approximations of the Pareto set. Lin et al. [151] propose learning a continuous approximation of the Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1). Subsequent improvements have been proposed to leverage diffusion models [152] and Stein variational gradient descent [153] for efficient solution generation, as well as incorporating data augmentation for Pareto set model building [154].
@@ 7.3 Computer Vision
The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has been successfully applied in autonomous driving [155, 156]. Since the encoder in computer vision usually contains large number of parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in some of the tasks. Hence, many methods are proposed to address the issue from the perspective of MOO, which use scalarization to balance multiple losses [157, 158, 159, 160, 161]. A comprehensive survey on multi-task dense prediction is in [4].
Besides the dense prediction task, MOO is also useful in other computer vision tasks such as point cloud [162, 163, 164], medical image denoising [165], and pose estimation [166].
@@ 7.4 Neural Architecture Search
Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently [167]. Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy. For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, and latency. These works are mainly based on gradient-based NAS methods (like DARTS [168]) and formulate it as a multi-objective optimization problem. Among them, [8, 169, 170, 171, 172] employ scalarization to identify a single solution, where the task and efficiency objectives are combined with fixed weights. However, altering the weights of the objectives needs a complete rerun of the search, which is computationally intensive. Therefore, Sukthanker et al. [173] propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
@@ 7.5 Recommendation System
Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty, diversity, serendipity, popularity, and others [174]. These factors naturally frame the problem as a multi-objective optimization optimization. Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives [175, 176, 177]. In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model [178, 179, 180, 181]. Additionally, some research efforts aim to identify finite Pareto sets [182] and infinite Pareto sets [183, 184], enabling the provision of personalized recommender models that cater to varying user preferences.
@@ 7.6 Large Language Models (LLMs)
Due to the powerful transferability of LLMs, users can fine-tune LLMs to specific downstream tasks or scenarios. This approach separates fine-tuning on each task, causing extensive costs in training and difficulties in deployment. MFTCoder [185] proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks. CoBa [186] studies multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects. MORLHF [187] and MODPO [188] train an LLM for every preference configuration by linearly combining multiple (implicit) reward models. To avoid retraining, some methods train multiple LLMs separately for each preference dimension and merge their parameters [189, 190] or output logits [191] to deal with different preference requirements. However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs. To improve efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [192, 193, 194] or model parameters [63].
@@ 7.7 Miscellaneous
In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [195, 196], federated learning [197, 198, 199], long-tailed learning [200, 201, 202], continual learning [203], diffusion models [204, 205, 206, 207], neural combinatorial optimization [208, 122, 209, 210, 211], GFlowNets [212, 213], and physics informed neural networks (PINNs) [214, 215]. These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
@@ 8 Resources
In this section, we introduce some benchmark datasets and open-source libraries for gradient-based MOO in deep learning. Benchmark datasets are summarized in Table III, with details in Appendix C.
Two popular libraries, LibMTL [^1] https://github.com/median-research-group/LibMTL [216] and LibMOON [^2] https://github.com/xzhang2523/libmoon [217], provide unified environments for implementing and fairly evaluating MOO methods. Both are built on PyTorch [218] and feature modular designs, enabling flexible development of new methods or application of existing ones to new scenarios. LibMTL [216] mainly focuses on finding a single Pareto-optimal solution and supports $24$ methods introduced in Section 3 and $6$ benchmark datasets. LibMOON [217] mainly focuses on exploring the whole Pareto set. It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5).
@@ 9 Challenges and Future Directions
Theoretical Understanding. While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
@@
Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received less attention. For instance, Chen et al. [24], as discussed in Section 6.2, provide the algorithm-dependent generalization analysis. Extending this to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
@@
For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets. Zhang et al. [232] established a generalization bound for HV-based Pareto set learning. However, the effect of various network architectures (Section 5.1) on approximating Pareto sets remain unclear. Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
@@
Reducing Gradient Balancing Costs. While gradient balancing methods in Section 3.2 often show strong performance in various applications, they come with significant computational overhead. Despite the introduction of some practical speedup strategies (discussed in Section 3.2.3), existing approaches remain insufficiently efficient. A deeper understanding of the optimization differences between gradient balancing, linear scalarization, and loss balancing is crucial. This insight could facilitate the integration of gradient balancing with linear scalarization and loss balancing, reducing computational overhead significantly and enabling its application in large-scale training scenarios.
@@
Dealing with Large Number of Objectives. Some real-world problems involve handling a large number of objectives, which poses significant challenges for current Pareto set learning algorithms in Section 5. As the number of objectives grows, the preference vector space expands exponentially, making it challenging for existing random sampling-based techniques to effectively learn the mapping between preference vectors and solutions. Developing efficient sampling strategies for high-dimensional preference spaces remains a important research problem. Additionally, exploring methods to reduce or merge objectives based on their properties can be a promising approach to minimize the total number of objectives.
@@
Distributed Training. Current MOO algorithms in deep learning are limited to single GPUs or machines, but scaling them to multi-GPU and distributed environments is increasingly important as models and datasets grow. This scaling introduces unique challenges compared to single-objective optimization. Gradient-balancing algorithms require efficient gradient distribution and communication across GPUs. Learning a set of Pareto-optimal solutions involves synchronizing multiple models across nodes with minimal overhead. Additionally, when data for different objectives is distributed across devices and cannot be shared (e.g., due to privacy constraints), collaborative computation of solutions or Pareto sets without direct data sharing becomes an important challenge.
@@
Advancements in Large Language Models (LLMs). As discussed in Section 7.6, current research in MOO for LLMs largely focuses on the RLHF stage. Expanding MOO techniques to other stages in the LLM lifecycle, such as pre-training and inference, is a valuable direction for future research. Addressing these challenges can lead to models that are better aligned with user needs across their entire development processes. Additionally, user preferences are currently modeled as a preference vector. However, this representation may oversimplify more complex user preferences on LLMs. Exploring advanced methods to capture and represent these intricate preferences can enhance the effectiveness and personalization of LLMs.
@@
Application in More Deep Learning Scenarios. While MOO has already been utilized in various deep learning scenarios, as highlighted in Section 7, there remain numerous unexplored areas within this field. Multi-objective characteristics are inherently present in most deep learning problems, as models are typically developed or evaluated based on multiple criteria. Consequently, trade-offs often arise naturally. Leveraging MOO methods to effectively address these trade-offs presents an opportunity for further exploration.